OpenStudy (idealist10):

Find a particular solution of y"-8y'+16y=23cosx-7sinx.

OpenStudy (idealist10):

@amistre64 @freckles @mathstudent55 @ParthKohli @uri

OpenStudy (amistre64):

what is your homogenous solution?

OpenStudy (idealist10):

The answer is \[y _{p}=\cos x-\sin x\]. But how do I do this problem?

OpenStudy (amistre64):

.... how i do it is ... i start by finding the =0 solution, the homogenous solution

OpenStudy (idealist10):

r=4, double root. y=e^(4x)(c1+c2x)

OpenStudy (amistre64):

now work a variation of parameters, let c1 and c2 be functions of x, and plug it into the equation

OpenStudy (idealist10):

I need to use a method involving the substitution of yp.

OpenStudy (amistre64):

let \[y_t=e^{4x}(f(x)+xg(x))\] what is the 1st and second derivatives?

OpenStudy (idealist10):

What's that ^ method above? Because I wanted to use the substitution \[y _{p}=Acos x-Bsin x\]

OpenStudy (amistre64):

variation of parameters ...

OpenStudy (amistre64):

your method looks like trial and error ... the guessing method

OpenStudy (amistre64):

16f e^(4x) + 16gx e^(4x) -32f e^(4x) -8ge^(4x) -32gx e^(4x) +(f'e^(4x)+ g'x e^(4x)) ( = 0 ) 4f' e^(4x) + 16f e^(4x) + g'e^(4x) + 4ge^(4x) + 4ge^(4x)+ 16gx e^(4x)+ 4g'x e^(4x) ---------------------------- 32f e^(4x) + 32gx e^(4x) -32f e^(4x) -8ge^(4x) -32gx e^(4x) +(f'e^(4x)+ g'x e^(4x)) ( = 0 ) 4f' e^(4x) + g'e^(4x) + 8ge^(4x) + 4g'x e^(4x) --------------------------- f'e^(4x)+ g'x e^(4x) = 0 f' 4e^(4x) + g'(e^(4x) + 4x e^(4x)) = 23cos(x)-7sin(x)

OpenStudy (jango_in_dtown):

@Idealist10 check the solution

OpenStudy (amistre64):

f'+ g'x = 0 4f' + g'(1+ 4x) = 23cos(x)-7sin(x)

OpenStudy (jango_in_dtown):

@Idealist10 did you check the attachment?

OpenStudy (idealist10):

Yes, I did. Thanks a lot for the help, guys.

OpenStudy (amistre64):

since f' = -g'x -4g'x + g'+ 4g'x = 23cos(x)-7sin(x) f' = -23x cos(x) + 7x sin(x) g' = 23cos(x) - 7sin(x) gx = 23xsin(x) + 7xcos(x) f = -23xsin(x) +7 sin(x)-7 x cos(x)-23 cos(x) f+gx = 7sin(x)-23cos(x)

OpenStudy (amistre64):

pfft, mine got off track someplace in there :)

OpenStudy (jango_in_dtown):

@amistre64 did you check my method?

OpenStudy (amistre64):

im not familiar enough with your method to comment critically about it

OpenStudy (amistre64):

im barely familiar with my own thougths :)

OpenStudy (jango_in_dtown):

Its shortcut methods for variation of parameters

OpenStudy (amistre64):

ive seen it, but i just do not have the practical experience with it

OpenStudy (jango_in_dtown):

Hmm we use it when we have sine or cosine functions as the tranformation to e^ix is easy// for example 1/f(D) e^ax is 1/f(a) e^ax if f(a) is not zero